OpenStudy (anonymous):

change this equation into slope-intercept form –x + 3y = 9

OpenStudy (nottim):

What is slope-interecept form, first of all (you can give teh defined one, or an example)

OpenStudy (anonymous):

rise over run

OpenStudy (nottim):

that's m=y/x right?

OpenStudy (lgbasallote):

didnt i just reply to that o.O

OpenStudy (nottim):

wait...so all oyu ahve to do is rearrange it?

OpenStudy (anikate):

3y=x+9

OpenStudy (anikate):

y=x/3+3

OpenStudy (anonymous):

i got x+9/3 :(

OpenStudy (nottim):

y=mx+b?

OpenStudy (anonymous):

yes

OpenStudy (nottim):

....uh